OpenStudy (anonymous):

Score: A rectangle has a length of x + 2, a width of 7, and a perimeter of 32. The value of x is

OpenStudy (jackellyn):

To find the perimeter, you add all the sides. So length x+2+x+2 and width 7+7 = 32 x+2+x+2+14=32 2x + 4 + 14 = 32 2x + 18 = 32 And solve for x (get x alone), do you know how to do that?